First National Bank pays 6.1% interest compounded semiannually. Second National Bank pays 6% interest compounded monthly. a. Calculate the effective annual rate for each bank. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

b. Which bank offers the higher effective annual rate?
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

First National Bank:

EAR=(1+0.061/2)^2-1

=6.19%(Approx)

Second National Bank:

EAR=(1+0.06/12)^12-1

=6.17%(Approx).

Hence  higher effective annual rate is for First National Bank.